Jonathan Macintosh is a highly successful Pennsylvania orchardman who has formed his own company to produce and package appleRequired 1 Required 2 Predict the energy cost for a month in which 87,000 pints of applesauce are produced. Energy cost
Answer #1

1) Variable cost per applesauce produced = Change in Cost/Change in unit = (73600-66400)/(136000-76000) = 0.12

Fixed cost = 73600-(136000*.12) = 57280

Total energy cost = 57280+0.12X

2) Total energy cost = 57280+0.12*87000 = $67720